# A non-vanishing conjecture for cotangent bundles on elliptic surfaces

### v2

@@ 1. Introduction
On the other hand, not much is known for cotangent bundles. Höring and Peternell [9] showed that Conjecture 1.1 partially holds for a smooth projective surface $S$ with $\kappa(S)\leq 1$, but the result remains open when $S$ admits an isotrivial elliptic fibration. Cao and Höring [5] proved that if a smooth projective variety $X$ admits an abelian fibration $f:X\rightarrow C$ onto a curve, then Conjecture 1.1 holds if $f_{\ast}\Omega_{X}^{1}$ has rank $1$, that is, $f$ is non-isotrivial. For a surface $S$ of general type, the Bogomolov vanishing theorem tells us that $\Omega_{S}^{1}$ is big if $c_{1}^{2}-c_{2}>0$, but little is known for the case $c_{1}^{2}-c_{2}\leq 0$.
@@
Notably, a (relatively minimal) isotrivial elliptic surface $f:S\rightarrow B$ with $\lambda^{f}(B)=0$ and that is not almost smooth has only one nontrivial symmetric differential (cf. [14, Section 4, Table III and (H)]).
In view of [9, Lemma A.1], one obtains the following consequence:
@@
Let $f:S\rightarrow B$ be a smooth elliptic surface. Then $\Omega_{S}^{1}$ is not pseudoeffective if and only if the fundamental group $\pi_{1}(S)$ is finite.
For the ‘if’ direction, suppose that $\Omega_{S}^{1}$ is pseudoeffective. We may assume that $f$ is relatively minimal and not almost smooth. Theorem 1.2, as well as [9, Proposition 5.4], states that either $g(B)\geq 1$ or $f$ has at least three multiple fibres. Thus, by [6, Theorem IV.9.12], there is a ramified covering $B^{\prime}\rightarrow B$ that induces a finite étale cover $S^{\prime}\rightarrow S$ such that the elliptic fibration $S^{\prime}\rightarrow B^{\prime}$ has no orbifold divisor. Now $\pi_{1}(S^{\prime})\simeq\pi_{1}(B^{\prime})$ is infinite as $g(B^{\prime})\geq 1$.
The proof of Theorem 1.2 basically uses the same idea as that of [9, Theorem 6.7]. However, to deal with singular fibres of types $II$, $III$ and $IV$, we need to work with a birational model $S^{\prime}$ of $S$ which might not be relatively minimal. As in [9, Corollary 6.12], we analyze the local obstruction for symmetric differentials on $S^{\prime}$ in terms of types of singular fibres. Then we use the fact from [9, Proposition 4.1] that the pseudoeffectivity of $\Omega_{S}^{1}$ is equivalent to that of $\Omega_{S^{\prime}}^{1}$, which completes the proof.
@@ 2. Elliptic surfaces
For a relatively minimal isotrivial elliptic fibration, the multiple singular fibres are not of type ${}_{m}I_{b}$ for $m\geq 2$ and $b\geq 1$; the non-multiple singular fibres are not of type $I_{b}$ or $I_{b}^{\ast}$ for $b\geq 1$.
@@
Choose a small disc $x\in\Delta\subset C$ such that $(C\times E)/G$ is locally isomorphic to $(\Delta\times E)/G_{x}$. Then the minimal resolution $S^{\prime}\rightarrow(C\times E)/G$ is locally isomorphic to that of $(\Delta\times E)/G_{x}$. If $G_{x}$ acts on $E$ by translation, the singularity is resolved by the multiple elliptic ${}_{m}I_{0}$ where $m=|G_{x}|$, so there is no $(-1)$-curve in the $f^{\prime}$-fibre under $x$. On the other hand, if $G_{x}$ fixes a point $e$ of $E$, then the exceptional divisor of the minimal resolution forms a Hirzebruch-Jung string by [2, Theorem III.5.4]. Indeed, for each nonzero element $g\in G_{x}$ and its fixed point $e^{\prime}\in E$, there is a local coordinate $(s,c)$ at $(x,e^{\prime})$ such that $g$ acts by $(s,c)\mapsto(\eta_{n}s,\eta_{n}^{\pm 1}c)$, where $n$ is the order of $g$ in $G_{x}$ and $\eta_{n}$ is a primitive $n$-th root of unity.
@@
Finally, when $G_{x}$ fixes a point of $E$, the corresponding $f^{\prime}$-fibre has a $(-1)$-curve if and only if the $f$-fibre under $x$ is of type $II$, $III$ or $IV$. In that case, the map $\mu$ is the composition of the following blowing down procedures. The curves contracted in each procedure are colored red.
@@ 3. The non-vanishing conjecture for cotangent bundles
Assume that $f^{\ast}\Omega_{B}^{1}(D)$ is not pseudoeffective, so that $B\simeq\mathbb{P}^{1}$. Under the setting (2.1), let $Z_{\ast}\subset C$ be the set of points under which fibres are singular of type $\ast\in\{I_{0}^{\ast}$, $II$, $II^{\ast}$, $III$, $III^{\ast}$, $IV$, $IV^{\ast}\}$. Let $e_{\ast}$ be the ramification index of the quotient map $\pi:C\rightarrow B$ at a point $x\in Z_{\ast}$, namely
@@ 3.1. Logarithmic symmetric differentials
First, we will address the singular fibres of type $I_{0}^{\ast}$, $II$, $III$ and $IV$. Consider $x\in Z_{III}$, and write $G_{x}=\langle g\rangle\simeq\mathbb{Z}/4\mathbb{Z}$. Choose a small disc $x\in\Delta\subset C$ such that $(C\times E)/G$ is locally isomorphic to $(\Delta\times E)/G_{x}$. Replacing the origin of $E$ if necessary, the fixed points of $g$ are $p_{0}=(x,[0])$ and $p_{1}=(x,[\frac{1}{2}+\frac{1}{2}i])$; the fixed point of $g^{2}$ is $p_{2}=(x,[\frac{1}{2}])$. Each quotient singularity of $(\Delta\times E)/G_{x}$ at $q_{\nu}=q(p_{\nu})$ is resolved by a rational curve $\Theta_{\nu}$. Let $\Theta\subset S^{\prime}|_{\Delta}$ be the proper transform of the central fibre of $(\Delta\times E)/G_{x}$ under $x$. Then the $f^{\prime}$-fibre under $x$ is $4\Theta+\Theta_{0}+\Theta_{1}+2\Theta_{2}$ whose intersection numbers are
@@
For $h_{\nu}<\frac{3m+1}{4}$ when $\nu=0,1$ and $h_{2}<\frac{m+1}{2}$, we have
(3.3) $H^{0}(S^{\prime}|_{\Delta}\setminus E_{0},\mathcal{B}_{\mathbf{h}})=H^{0}(S^{\prime}|_{\Delta}\setminus E_{0},S^{m}\Omega_{S^{\prime}}^{1})\simeq H^{0}(S^{\prime}|_{\Delta},\mathcal{B}_{\mathbf{h}}).$
This result does not hold for larger ones.
@@
(3.4) $\begin{split}H^{0}(\Delta\times E,S^{m}\Omega_{\Delta\times E}^{1})^{G_{x}}&\simeq H^{0}(\Delta\times E\setminus\{p_{0},p_{1},p_{2}\},S^{m}\Omega_{\Delta\times E}^{1})^{G_{x}}\\ &\simeq H^{0}(S^{\prime}|_{\Delta}\setminus E_{0},S^{m}\Omega_{S^{\prime}}^{1}),\end{split}$
which is true as $\Delta\times E$ is smooth and $S^{m}\Omega_{\Delta\times E}^{1}$ is reflexive, one can identify the space of logarithmic symmetric differentials on $S^{\prime}$ as a subspace of symmetric differentials on $\Delta\times E$.
@@
Choose local coordinates $(s,c_{\nu})$ at $p_{\nu}$ such that $g:(s,c_{\nu})\mapsto(\eta_{4}s,\eta_{4}c_{\nu})$ for $\nu=0,1$ and $g^{2}:(s,c_{2})\mapsto(-s,-c_{2})$. Then by [12, II, p.583] there exist coordinate charts $W_{\nu 1}$ and $W_{\nu 2}$ whose respective local coordinates are $(y_{\nu},s_{\nu})$ and $(x_{\nu},t_{\nu})$ such that
@@
In these charts, $\Theta_{\nu}$ is defined by $y_{\nu}=x_{\nu}=0$ and $\Theta$ is defined by $s_{\nu}=0$. Moreover, $\Theta$ does not meet $W_{\nu 2}$.
For $\nu=0,1$, the map (3.4) is locally written on $W_{\nu 0}$ as
@@
$ds\mapsto 4^{-1}y_{\nu}^{-3/4}(4y_{\nu}ds_{\nu}+s_{\nu}dy_{\nu}),\qquad dc_{\nu}\mapsto 4^{-1}y_{\nu}^{-3/4}dy_{\nu};$
on $W_{\nu 1}$, the map
@@
$ds\mapsto 4^{-1}x_{\nu}^{-3/4}dx_{\nu},\qquad dc_{\nu}\mapsto 4^{-1}x_{\nu}^{-3/4}(4x_{\nu}dt_{\nu}+t_{\nu}dx_{\nu}).$
Hence we get a valuation $\mathrm{ord}_{\Theta_{\nu}}$ with
@@
one can see that $\frac{dy_{\nu}}{y_{\nu}}$ does not define a logarithmic symmetric differential on $W_{\nu 1}\cup W_{\nu 2}$, whilst $y_{\nu}(\frac{dy_{\nu}}{y_{\nu}})^{4}$ does. Similarly, $\frac{dy_{2}}{y_{2}}$ does not define a logarithmic symmetric differential on $W_{21}\cup W_{22}$, while $y_{2}(\frac{dy_{2}}{y_{2}})^{2}$ does. In this fashion, one can see that
@@
    If $\ast=I_{0}^{\ast}$, the $f^{\prime}$-fibre under $x$ is $2\Theta+\Theta_{0}+\Theta_{1}+\Theta_{2}+\Theta_{3}$ with self-intersection $-2$ and $(\Theta.\Theta_{\nu})=1$. For a tuple $\mathbf{h}\in\mathbf{Z}_{\geq 0}^{4}$, we have
@@
    if and only if $h_{\nu}<\frac{m+1}{2}$, where $E_{0}=\sum_{\nu}\Theta_{\nu}$.
@@
    If $\ast=II$, the $f^{\prime}$-fibre under $x$ is $6\Theta+\Theta_{0}+2\Theta_{1}+3\Theta_{2}$ with self-intersection $(\Theta^{2})=-1$, $(\Theta_{\nu}^{2})=-6/(\nu+1)$ and $(\Theta.\Theta_{\nu})=1$. For a tuple $\mathbf{h}\in\mathbf{Z}_{\geq 0}^{3}$, we have
@@
    if and only if $h_{\nu}<\frac{(6/(\nu+1)-1)m+1}{6/(\nu+1)}$, where $E_{0}=\sum_{\nu}\Theta_{\nu}$.
@@
    If $\ast=IV$, the $f^{\prime}$-fibre under $x$ is $3\Theta+\Theta_{0}+\Theta_{1}+\Theta_{2}$ with self-intersection $(\Theta^{2})=-1$, $(\Theta_{\nu}^{2})=-3$ and $(\Theta.\Theta_{\nu})=1$. For a tuple $\mathbf{h}\in\mathbf{Z}_{\geq 0}^{3}$, we have
@@
    if and only if $h_{\nu}<\frac{2m+1}{3}$, where $E_{0}=\sum_{\nu}\Theta_{\nu}$.

### v3

@@ 1. Introduction
On the other hand, not much is known for cotangent bundles. Höring and Peternell [9] showed that Conjecture 1.1 partially holds for a smooth projective surface $S$ with $\kappa(S)\leq 1$, but the result remains open when $S$ admits an isotrivial elliptic fibration. Cao and Höring [5] proved that if a smooth projective variety $X$ admits an abelian fibration $f:X\rightarrow C$ onto a curve, then Conjecture 1.1 holds if $f_{\ast}\Omega_{X}^{1}$ has rank $1$, which corresponds to the non-isotrivial one for elliptic surfaces. For a surface $S$ of general type, the Bogomolov vanishing theorem tells us that $\Omega_{S}^{1}$ is big if $c_{1}^{2}-c_{2}>0$, but little is known for the case $c_{1}^{2}-c_{2}\leq 0$.
@@
Theorem 1.2.
Let $f:S\rightarrow B$ be a relatively minimal isotrivial elliptic surface. If $\Omega_{S}^{1}$ is pseudoeffective, then $H^{0}(S,S^{m}\Omega_{S}^{1})\neq 0$ for some $m>0$.
Our main theorem can be proven by applying [9, Proposition 5.2 and 4.6] to the following technical lemma:
@@
For a given relatively minimal elliptic surface $f:S\rightarrow B$, let $\lambda^{f}(B)$ be the Iitaka dimension of the $\mathbb{Q}$-line bundle $K_{B}+\sum_{i=1}^{s}(1-\frac{1}{\nu_{i}})a_{i}$, where $f$ has multiple fibres over $a_{i}\in B$ with multiplicity $\nu_{i}$. Notably, a (relatively minimal) isotrivial elliptic surface $f:S\rightarrow B$ with $\lambda^{f}(B)=0$ and that is not almost smooth has only one nontrivial symmetric differential (cf. [14, Section 4, Table III and (H)]).
In view of [9, Lemma A.1], one can characterize the pseudoeffectivity of the cotangent bundle on an elliptic surface:
@@
Let $f:S\rightarrow B$ be a smooth elliptic surface. Then $\Omega_{S}^{1}$ is pseudoeffective if and only if the fundamental group $\pi_{1}(S)$ is infinite.
The ‘if’ direction is due to [9, Lemma A.1]. For the converse, suppose that $\Omega_{S}^{1}$ is pseudoeffective. We may assume that $f$ is relatively minimal and not almost smooth. Theorem 1.3, as well as [9, Proposition 5.4], states that either $g(B)\geq 1$ or $f$ has at least three multiple fibres. Thus, by [6, Theorem IV.9.12], there is a ramified covering $B^{\prime}\rightarrow B$ that induces a finite étale cover $S^{\prime}\rightarrow S$ such that the elliptic fibration $S^{\prime}\rightarrow B^{\prime}$ has no orbifold divisor. Now $\pi_{1}(S^{\prime})\simeq\pi_{1}(B^{\prime})$ is infinite as $g(B^{\prime})\geq 1$.
The proof of Theorem 1.3 basically uses the same idea as that of [9, Theorem 6.7]. However, to deal with singular fibres of types $II$, $III$ and $IV$, we need to work with a birational model $S^{\prime}$ of $S$ which might not be relatively minimal. Using an explicit computation, we analyze the local obstruction for symmetric differentials on $S^{\prime}$ in terms of types of singular fibres. Then we use the fact from [9, Proposition 4.1] that the pseudoeffectivity of $\Omega_{S}^{1}$ is equivalent to that of $\Omega_{S^{\prime}}^{1}$, which completes the proof.
@@ 2. Elliptic surfaces
For a relatively minimal isotrivial elliptic fibration, the multiple singular fibres are of type ${}_{m}I_{0}$ for $m\geq 2$; the non-multiple singular fibres are of type $I_{0}^{\ast}$, $II$, $II^{\ast}$, $III$, $III^{\ast}$, $IV$ or $IV^{\ast}$.
@@
Choose a small disc $x\in\Delta\subset C$ such that $(C\times E)/G$ is locally isomorphic to $(\Delta\times E)/G_{x}$. Then the minimal resolution $S^{\prime}\rightarrow(C\times E)/G$ is locally isomorphic to that of $(\Delta\times E)/G_{x}$. If $G_{x}$ acts on $E$ by translation, the $f^{\prime}$-fibre under $x$ is the multiple elliptic ${}_{m}I_{0}$ where $m=|G_{x}|$, so there is no $(-1)$-curve in the fibre. On the other hand, if $G_{x}$ fixes a point $e$ of $E$, then the exceptional divisor of the minimal resolution forms a Hirzebruch-Jung string by [2, Theorem III.5.4]. Indeed, for each nonzero element $g\in G_{x}$ and its fixed point $e^{\prime}\in E$, there is a local coordinate $(s,c)$ at $(x,e^{\prime})$ such that $g$ acts by $(s,c)\mapsto(\eta_{n}s,\eta_{n}^{\pm 1}c)$, where $n$ is the order of $g$ in $G_{x}$ and $\eta_{n}$ is a primitive $n$-th root of unity.
@@
Table 2.1. Monodromy groups and quotient singularities according to the type of singular fibres.
@@
Finally, when $G_{x}$ fixes a point of $E$, the corresponding $f^{\prime}$-fibre has a $(-1)$-curve if and only if the $f$-fibre under $x$ is of type $II$, $III$ or $IV$. In that case, the map $\mu$ is the composition of the following blowing down procedures.
Figure 2.1. Blowing-down procedures. The notation $a(-b)$ indicates that the corresponding curve has multiplicity $a$ and self-intersection $-b$. The curves contracted in each procedure are colored red.
@@ 3. The non-vanishing conjecture for cotangent bundles
Assume that $f^{\ast}\Omega_{B}^{1}(D)$ is not pseudoeffective, so that $B\simeq\mathbb{P}^{1}$ and $\sum_{i=1}^{s}(1-\frac{1}{\nu_{i}})<2$. Under the setting (2.1), let $Z_{\ast}\subset C$ be the set of points under which fibres are singular of type $\ast\in\{I_{0}^{\ast}$, $II$, $II^{\ast}$, $III$, $III^{\ast}$, $IV$, $IV^{\ast}\}$. Let $e_{\ast}$ be the ramification index of the quotient map $\pi:C\rightarrow B$ at a point $x\in Z_{\ast}$, namely
@@ 3.1. Logarithmic symmetric differentials
First, we will address the singular fibres of type $I_{0}^{\ast}$, $II$, $III$ and $IV$. Consider $x\in Z_{III}$, and write $G_{x}=\langle g\rangle\simeq\mathbb{Z}/4\mathbb{Z}$. Choose a small disc $x\in\Delta\subset C$ such that $(C\times E)/G$ is locally isomorphic to $(\Delta\times E)/G_{x}$. Replacing the origin of $E$ if necessary, the fixed points of $g$ are $p_{0}=(x,[0])$ and $p_{1}=(x,[\frac{1}{2}+\frac{1}{2}i])$; the fixed point of $g^{2}$ is $p_{2}=(x,[\frac{1}{2}])$. The quotient singularities of $(\Delta\times E)/G_{x}$ at $q_{\nu}=q(p_{\nu})$ for $\nu=0,1$ are $A_{4,1}$-singularities, so they are resolved by $(-4)$-curves $\Theta_{\nu}$; that of $q_{2}=q(p_{2})$ is an $A_{1}$-singularity, so it is resolved by a $(-2)$-curve $\Theta_{2}$. Let $\Theta\subset S^{\prime}|_{\Delta}$ be the proper transform of the central fibre of $(\Delta\times E)/G_{x}$ under $x$. As shown by Figure 2.1, the $f^{\prime}$-fibre under $x$ is $4\Theta+\Theta_{0}+\Theta_{1}+2\Theta_{2}$ whose intersection numbers are
@@
For $\mathbf{h}\in\mathbb{Z}_{\geq 0}^{3}$, we have
(3.3) $H^{0}(S^{\prime}|_{\Delta}\setminus E_{0},\mathcal{B}_{\mathbf{h}})=H^{0}(S^{\prime}|_{\Delta}\setminus E_{0},S^{m}\Omega_{S^{\prime}}^{1})\simeq H^{0}(S^{\prime}|_{\Delta},\mathcal{B}_{\mathbf{h}})$
if and only if $h_{\nu}<\frac{3m+1}{4}$ for $\nu=0,1$ and $h_{2}<\frac{m+1}{2}$.
@@
(3.4) $\begin{split}H^{0}(\Delta\times E,S^{m}\Omega_{C\times E}^{1})^{G_{x}}&\simeq H^{0}(\Delta\times E\setminus\{p_{0},p_{1},p_{2}\},S^{m}\Omega_{C\times E}^{1})^{G_{x}}\\ &\simeq H^{0}(S^{\prime}|_{\Delta}\setminus E_{0},S^{m}\Omega_{S^{\prime}}^{1}),\end{split}$
which is true as $\Delta\times E$ is smooth and $S^{m}\Omega_{C\times E}^{1}$ is reflexive, one can identify the space of logarithmic symmetric differentials on $S^{\prime}$ as a subspace of symmetric differentials on $\Delta\times E$.
@@
Choose neighborhoods $\Delta_{\nu}$ of $p_{\nu}$ with respective local coordinates $(s,c_{\nu})$ such that $g:(s,c_{\nu})\mapsto(\eta_{4}s,\eta_{4}c_{\nu})$ for $\nu=0,1$ and $g^{2}:(s,c_{2})\mapsto(-s,-c_{2})$. Then by [12, II, p.583] there exist coordinate charts $W_{\nu 1}$ and $W_{\nu 2}$ on $S^{\prime}|_{\Delta}$ whose respective local coordinates are $(y_{\nu},s_{\nu})$ and $(x_{\nu},t_{\nu})$ such that
@@
In those charts, $\Theta_{\nu}$ is defined by $y_{\nu}=x_{\nu}=0$ and $\Theta$ is defined by $s_{\nu}=0$. Moreover, $\Theta$ does not meet $W_{\nu 2}$. We visualize the setting in Figure 3.1.
Figure 3.1. Coordinate charts on the $f^{\prime}$-fibre of type $III$.
For $\nu=0,1$, following [3, Section 3.2], one can describe the valuation $\mathrm{ord}_{\Theta_{\nu}}$ on
$\bigoplus_{m\in\mathbb{Z}_{\geq 0}}H^{0}(\Delta_{\nu},S^{m}\Omega_{C\times E}^{1})\simeq\mathbb{C}[s,c_{\nu},ds,dc_{\nu}]$
as follows. It is expressed using the ring map
@@
$ds\mapsto 4^{-1}y_{\nu}^{-3/4}(4y_{\nu}ds_{\nu}+s_{\nu}dy_{\nu}),\qquad dc_{\nu}\mapsto 4^{-1}y_{\nu}^{-3/4}dy_{\nu}$
on $W_{\nu 1}$, and using the ring map
@@
$ds\mapsto 4^{-1}x_{\nu}^{-3/4}dx_{\nu},\qquad dc_{\nu}\mapsto 4^{-1}x_{\nu}^{-3/4}(4x_{\nu}dt_{\nu}+t_{\nu}dx_{\nu})$
on $W_{\nu 2}$. Hence the valuation $\mathrm{ord}_{\Theta_{\nu}}$ satisfies
@@
one can see that $\frac{dy_{\nu}}{y_{\nu}}$ does not define a logarithmic symmetric differential on $W_{\nu 1}\cup W_{\nu 2}$, whilst $y_{\nu}(\frac{dy_{\nu}}{y_{\nu}})^{4}$ does.
Similarly, for $\nu=2$, the valuation $\mathrm{ord}_{\Theta_{2}}$ is expressed using
$\mathbb{C}[s,c_{2},ds,dc_{2}]\rightarrow\mathbb{C}(y_{2}^{1/2})[s_{2},dy_{2},ds_{2}]$
given by
$s\mapsto y_{2}^{1/2}s_{2},\qquad c_{2}\mapsto y_{2}^{1/2},$
$ds\mapsto 2^{-1}y_{2}^{-1/2}(2y_{2}ds_{2}+s_{2}dy_{2}),\qquad dc_{2}\mapsto 2^{-1}y_{2}^{-1/2}dy_{2}$
on $W_{21}$, and using the ring map
$\mathbb{C}[s,c_{2},ds,dc_{2}]\rightarrow\mathbb{C}(x_{2}^{1/2})[t_{2},dx_{2},dt_{2}]$
given by
$s\mapsto x_{2}^{1/2},\qquad c_{2}\mapsto x_{2}^{1/2}t_{2},$
$ds\mapsto 2^{-1}x_{2}^{-1/2}dx_{2},\qquad dc_{2}\mapsto 2^{-1}x_{2}^{-1/2}(2x_{2}dt_{2}+t_{2}dx_{2})$
on $W_{22}$. Thus one can see that $\frac{dy_{2}}{y_{2}}$ does not define a logarithmic symmetric differential on $W_{21}\cup W_{22}$, while $y_{2}(\frac{dy_{2}}{y_{2}})^{2}$ does. In this fashion, one concludes that
@@
    If $\ast=I_{0}^{\ast}$, the $f^{\prime}$-fibre under $x$ is $2\Theta+\Theta_{0}+\Theta_{1}+\Theta_{2}+\Theta_{3}$ with self-intersection $-2$ and $(\Theta.\Theta_{\nu})=1$ for all $\nu$. For a tuple $\mathbf{h}\in\mathbf{Z}_{\geq 0}^{4}$, we have
@@
    if and only if $h_{\nu}<\frac{m+1}{2}$ for all $\nu$, where $E_{0}=\sum_{\nu}\Theta_{\nu}$.
@@
    If $\ast=II$, the $f^{\prime}$-fibre under $x$ is $6\Theta+\Theta_{0}+2\Theta_{1}+3\Theta_{2}$ with self-intersection $(\Theta^{2})=-1$, $(\Theta_{\nu}^{2})=-6/(\nu+1)$ and $(\Theta.\Theta_{\nu})=1$ for all $\nu$. For a tuple $\mathbf{h}\in\mathbf{Z}_{\geq 0}^{3}$, we have
@@
    if and only if $h_{\nu}<\frac{(6/(\nu+1)-1)m+1}{6/(\nu+1)}$ for all $\nu$, where $E_{0}=\sum_{\nu}\Theta_{\nu}$.
@@
    If $\ast=IV$, the $f^{\prime}$-fibre under $x$ is $3\Theta+\Theta_{0}+\Theta_{1}+\Theta_{2}$ with self-intersection $(\Theta^{2})=-1$, $(\Theta_{\nu}^{2})=-3$ and $(\Theta.\Theta_{\nu})=1$ for all $\nu$. For a tuple $\mathbf{h}\in\mathbf{Z}_{\geq 0}^{3}$, we have
@@
    if and only if $h_{\nu}<\frac{2m+1}{3}$ for all $\nu$, where $E_{0}=\sum_{\nu}\Theta_{\nu}$.